Give an example of a continuously differentiable function from R2 to \mathbb{R} , which has an isolated local maximum at (0,0) and in (-17,9) and (0,3) an isolated local minimum in each case. Justify your answer.
Answer #1

A function having local maxima at (0,0) and local minimum at (-17,9) is

f(x,y)=\left [1-\left (\frac{ x^2}{17^2} +\frac{y^2}{9^2} \right ) \right ]^2

which possesses local maximum value equal to 1 at (0,0) and local minimum value 0 at (-17,9) in the subset

D=\{(x,y)\in\mathbb R^2: \frac{x^2}{17^2}+\frac{y^2}{9^2}\le 1\}

Similarly, for local maximum at (0,0) and local minimum at (0,3) an example is

g(x,y)=\left [ 1-\left (x^2+\frac{y^2}{3^2}\right )\right ]^2

which attains local maximum value equal to 1 at (0,0) and local minimum value equal to 0 at (0,3) in the subset

E=\{(x,y)\in\mathbb R^2:x^2+\frac{y^2}{3^2}\le 1\}